In the year 2015, Anna bought a new car for $36,000. In 2017, she was told that the value of her car was $25,000 due to depreciation. She is told that the value of her car depreciates linearly.
1. Find a function V(t) which gives the value of the car t years after the year 2015.
2. In 2018, Anna is told that she will be given $15,000 for her car if she decides to trade it in for a new car. Use the function from part (1) above to determine the value of her car in 2018.
3. Is the $15,000 value fair based on what she was told about linear depreciation? Explain your answer?

Answer by Boreal(15235)

Let 2015 be t=0, 2016 be t=1, ...
It depreciated $11000 in 2 years, or $5500 a year
V(t)=36000-5500t function
2018 will be t=3
V(3)=36000-5500*3=$19500
It would appear not to be fair based on depreciation, especially since in 2018 can be considered year 3 in January or December. The problem isn't clear on when in the year the clock starts or stops
Since cars don't depreciate linearly but a great deal at first, if she were told this was a linear issue, she was not told the right thing.